15. November 2007. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB Nr. Institutsbibliothek Nr., LB = HB–Lehrbuchsammlung, LS = HB–Lesesaal

Literatur zum Problemlösen

[1] Problemlösen lernen, mathematik lehren, Heft 52. Friedrich, 1992. ISSN 0175-2235.

[2] Art of Problem Solving. Art of Problem Solving. http://www.artofproblemsolving.com, Gesehen Oktober 2006. Hinweis in CMJ 37:4 (2006), 329 unten. Siehe auch [24, 25, 56].

[3] Ludwig Bauer. Mathematische Fhigkeiten: mathematische Fhigkeiten in der Sekundarstufe II und ihre Bedeutung fr das Lsen von Abituraufgaben. Uni-Taschenbcher 835: Didaktik der Mathematik. Paderborn: Schningh, 1978. HB: BB1435. Vollstndig zugl.: Regensburg, Univ., Diss., 1977.

[4] Ludwig Bauer. Pyramiden. Anregungen zu einem problemorientierten Mathematikunterricht. In Peter Baptist, editor, Mathematikunterricht im Wandel. Bausteine für den Unterricht, pages 69–97. C. C. Buchner, 2000. HB: Kb919.

[5] Vasile Berinde. Exploring, Investigating and Discovering in Mathematics. Birkhäuser, 2004. ISBN 3-7643-7019-X. HBZ. A concrete and as exact as possible description of the act of discovery in mathematics (p. xv). The target audience are mainly highschool teachers, .., and teachers in mathematics. Review: http://www.maa.org/reviews/exploringinvestigating.html.

[6] William Biggs. Ants, Bikes, and Clocks: Problem Solving for Undergraduates. SIAM, 2004. ISBN 0-89871-574-1. HBZ.

[7] N. A. Branca. Mathematical problem solving: lessons from the British experience. In Edward Allen Silver, editor, Teaching and Learning Mathematical Problem Solving: Multiple Research Perspectives (San Diego 1983), pages 71–80. Erlbaum, 1985. Siehe [47].

[8] Regina Bruder. Entdeckend und problemlösend Mathematik lernen – aber wie? Vortrag Aachen 24.06.2003: Einladung und ausgedruckte Folien, 2003.

[9] Gengzhe Chang and Thomas W. Sederberg. Over and Over Again. The New Mathematical Library. MAA, 1997. ISBN 0-88385-641-7. Suitable as supplemental reading in courses in differential and integral calculus, numerical analysis, approximation theory and computer-aided geometric design. Ed Sandifer: The real contents of the book is contest problem solving and how it is relasted to the underlying mathematics.

[10] Ulrich Daepp and Pamela Gorkin. Reading, writing, and proving. A closer look at mathematics. Untergraduate Texts in Mathematics. New York: Springer, 2003.

[11] J. P. D’Angelo and D. B. West. Mathematical Thinking: Problem-Solving and Proofs. Prentice Hall, 1997. ISBN 0-13-263393-0. MB: 18192. Einfhrung / Analysis / Diskrete Mathematik / Elementare Zahlentheorie.

[12] D. Dörner. Die kognitive Organisation beim Problemlösen: Versuche zu einer kybernetischen Theorie der elementaren Informationsverarbeitungsprozesse beim Denken. Stuttgart: Huber, 1974. ISBN 3-456-30595-8. HB: Aa7167 (auch LB). Inst. f. Psychologie mit Schwerpunkt Berufliche Rehabilitation, AC, Jägerstr. zw. Nr. 17 und 19, Erdgeschoss, Zi. 004: AL.730.DÖR(2).

[13] D. Dörner. Problemlösen als Informationsverarbeitung. Stuttgart: Kohlhammer, 1976, ^{2}1979, 1987. ISBN 3-17-009711-3. HB: Aa9606. Inst. f. Psychologie mit Schwerpunkt Berufliche Rehabilitation, AC, Jägerstr. zw. Nr. 17 und 19, Erdgeschoss, Zi. 004: AL.730.DÖR.

[14] Arthur Engel. Problem-Solving Strategies. Problem Books in Mathematics. Springer-Verlag, 1998. ISBN 0-387-98219-1, HBZ. Aufgabensammlung.

[15] J. Douglas Faires. First Steps for Math Olympians. Using the American Mathematics Competitions. Problem Books. MAA, 2006. ISBN 0-88385-824-X. The AMC have been given for more than fifty years to millions of high school students. This book considers the basic ideas behind the solutions to the majority of these problems.

[16] N. Fredericksen. Implications of cognitive theory for instruction in problem solving. Review of Educational Research, 54:363–407, 1984. HB: Z1865.

[17] J. W. Getzels. Creative thinking, problem solving, and instruction. In Ernest R. Hilgard, editor, Theories of Learning and Instruction (63rd Yearbook of the National Society for the Study of Education, Part I), pages 240–267. Univ. of Chicago Press, 1964. Deutsch in: Ernest R. Hilgard (bzw. Gordon H. Bower), Theorien des Lernens, Band 1, 1970, ^{5}1984 und Band 2, 1971, ^{3}1984 (HB: Aa6532*).

[18] Nicola Haas. Das Extremalprinzip als Element mathematischer Denk- und Problemlöseprozesse. Diss. Aachen, 1999.

[19] Frank Heinrich and Bernd Zimmermann, editors. Problemlösen, Der Mathematikunterricht 49, Heft 1. Seelze: Friedrich, 2003. Enthält: S. 5–8: Hartmut Rehlich, Ideen zur Organisation entdeckenden Lernens im Mathematikunterricht; S. 19–24: Michael Schmitz: Verwendung eines CAS bei einer geometrischen Problemstellung im Rahmen einer Seminarfacharbeit; S. 25–41: Torsten Fritzlar: Zur Sensibilisierung von Studierenden für die Komplexität problemorientierten Mathematikunterrichts – eine Herausforderung für die Lehrerausbildung (?); S. 42–57: Bernd Zimmermann, Mathematisches Problemlösen und Heuristik in einem Schulbuch; S. 58–69: Lenni Haapasalo, Von der Katastrophe zum Erfolg mit den Brüchen - Ein möglicher Weg zur Lösung von Problemen der Bruchrechnung.

[20] H.-W. Henn. Computergestütztes problemlösen in der geometrie. Der Mathematikunterricht, 40(1):25–38, 1994. HB: Z 5577. Heftthema([?]): Computereinsatz im Geometrieunterricht der Sekundarstufe I (Henn).

[21] Michael B. Jackson and John R. Ramsay, editors. Problems for Student Investigation, Textbook Supplement. MAA, 1993. ISBN 0-88385-086-9. HBZ. How to use calculus to solve real problems, how to use the library .., how to read and write mathematical material, and how to cooperate with their peers ..

[22] Charles A. Jones. A logo-based course in problem solving. In D. A. Smith, G. J. Porter, L. C. Leinbach, and R. H. Wenger, editors, Computers and Mathematics. The Use of Computers in Undergraduate Instruction, MAA Notes 9, pages 97–100. MAA, 1988. FL.

[23] S. G. Krantz. Techniques of Problem Solving. AMS, 1996. ISBN 0-8218-0619-X. IB 319: Ba008. FL: UB Duisburg 01 TBG 1986.

[24] Sandor Lehoczky and Richard Rusczyk. The Art of Problem Solving Volume 1: the Basics, publisher = Stanford, Calif.: Greater Testing Concepts, series = , year = 1993, note = HBZ. Volume 1 is appropriate for students just beginning in math contests. MATHCOUNTS and novice high school students particularly have found it invaluable. Although the Art of Problem Solving is widely used by students preparing for mathematics competitions, the book is not just a collection of tricks. The emphasis on learning and understanding methods rather than memorizing formulas enables students to solve large classes of problems beyond those presented in the book ,.

[25] Sandor Lehoczky and Richard Rusczyk. The Art of Problem Solving Volume 2: and Beyond. Stanford, Calif.: Greater Testing Concepts, 1994. HBZ.

[26] Manfred Müller. Mathematisches Denken – Ein hochschuldidaktischer Beitrag zu den Wissenschaftsstrukturen der Mathematik aus erziehungsphilosophischer Sicht. Erziehungsphilosophie 6. Frankfurt am Main: Verlag Peter Lang, 1985. ISBN 3-8204-5649-X. Habil.-Schr. Bonn 1985. Will einen konkreten Beitrag zur Verbesserung der Ausbildung von Studierenden der Mathematik un der Didaktik der Mathematik leisten (S. 1). Reflexion über Mathematik und mathemaische Erkenntnis. Beispiele für grundlegende mathematische Methoden. HB: KB1549-6+1. Review: PM 31:3 (1989), 189–190.

[27] Steve Olson. Count down: six kids vie for glory at the world’s toughest math competition. Boston [u.a.]: Houghton Mifflin, 2004. ISBN 0-618-25141-3. HBZ. Review: CMJ 37:4 (2006), 328–331 (sehr empfehlend).

[28] Karlheinz Petzel. Die Quadratur krummliniger Figuren oder: Anleitung zur Lösung wissenschaftlicher Probleme. Stuttgart: Hirzel, 1999. ISBN 3-7776-0865-3. HBZ.

[29] G. Pólya. Vom Lösen mathematischer Aufgaben: Band 2: publisher = Birkhäuser, series = , year = 1967, note = MB: 3865. Band 1: [36] .

[30] G. Pólya. Schule des Denkens: vom Lösen mathematischer Probleme. Sammlung Dalp 36. Bern: Francke, 1949, ^{4}1995. ISBN 3-7720-0608-6. Engl. Übersetzung ”How to Solve It“ [33]. HBZ. FH (Eupener Str.): TBP 137(4).

[31] G. Pólya. Mathematics and Plausibel Reasoning. Vol. 1: Induction and Analogy in Mathematics. Princeton Univ. Press, 1954. MB: 896a.

[32] G. Pólya. Mathematics and Plausibel Reasoning. Vol. 2: Patterns of Plausible Inference. Princeton Univ. Press, 1954. MB: 896b.

[33] G. Pólya. How to Solve It: a New Aspect of Mathematical Method. Penguin mathematics. New York: Penguin Books, 1957 (Princeton U. Press), ^{2}1990. ISBN 0-14-012499-3. HBZ: 361 BI. Übersetzung von [30].

[34] G. Pólya. Mathematik und plausibles Schließen. Band 1: Induktion und Analogie in der Mathematik. Wissenschaft und Kultur 14. Birkhäuser, 1962, ^{3}1988. MB: 2308, 2395. Originalausgabe: Mathematics and Plausible Reasoning. Vol. 1: Induction and Analogy in Mathematics, Princeton Univ. Press, 1954.

[35] G. Pólya. Mathematik und plausibles Schließen. Band 2: Typen und Strukturen plausibler Folgerung. Wissenschaft und Kultur 15. Birkhäuser, 1963. MB: 2606. Originalausgabe: Mathematics and Plausible Reasoning. Vol. 2: Patterns of Plausible Inference, Princeton Univ. Press, 1954.

[36] G. Pólya. Vom Lösen mathematischer Aufgaben: Band 1: Einsicht und Entdeckung, Lernen und Lehren. Birkhäuser, 1966. MB: 3660. Band 2: [29].

[37] G. Pólya. Wie lehren wir Problemlösen? Mathematiklehrer, 1980(1):3–5, 1980. ISSN 0720-0420.

[38] Dave Pratt. Computer modelling in a problem-solving environment. In David Pimm, editor, Mathematics, Teachers and Children. A Reader, pages 59–67. London: Hodder & Stoughton, 1988. ISBN 0-340–48756-9. HBZ. Digger and lorries with waiting probabilities. Ein gutes Beispiel zum Modellieren. Offene Diskussion.

[39] Alan H. Schoenfeld. Mathematical Problem Solving. Acad. Press, 1985, ^{2}1987, ^{3}1988. ISBN 0-12-628870-4. HBZ: 361, 466, 468, 61, 708; 2. Aufl.: 464; 3. Aufl.: 290. FL: UB Bielefeld 100/395683+01. What is ”mathematical thinking“? 56–57: van Hiele’s five levels of thought in geometry: 1 Recognition, 2 Analysis, 3 Ordering, 4 Deduction, 5 Rigor. 76–96: Toward more precise and usable descriptions of heuristic strategies (exa. ”examining special cases“ and ”establishing subgoals“). 108–117: Control, frequently used heuristics. 388–397: References.

[40] Alan H. Schoenfeld, editor. Cognitive Science and Mathematics Education. Erlbaum, 1987. HBZ: 290, 361, 385, 464, 465, 466, 467, 708. ISBN 0-89859-791-9.

[41] Alan H. Schoenfeld. Teaching mathematical thinking and problem solving. In Lauren B. Resnick and Leopold E. Klopfer, editors, Toward the Thinking Curriculum: Current Cognitive Research, Yearbook of the Association for Supervision and Curriculum Development, pages 83–103. Association for Supervision and Curriculum Development, 1989. FL: UB Bielefeld 164/1332060+1.

[42] Alan H. Schoenfeld. On mathematics as sense-making: an informal attack on the unfortumate divorce of formal and informal mathematics. In D. Perkins, J. Segal, and J. Voss, editors, Informal Reasoning and Education. Erlbaum, 1989 oder 1990.

[43] Alan H. Schoenfeld, editor. A Source Book for College Mathematics Teaching: A Report from the MAA Committee on the Teaching of Undergaduate Mathematics. MAA, 1990. HBZ: 290, 361. ISBN 0-88385-068-0.

[44] Alan H. Schoenfeld, editor. Mathematical Thinking and Problem Solving, Studies in mathematical thinking and learning. Erlbaum, 1994. HBZ: 61, 361, 385, 464, Kob7. ISBN 0-8058-0990-2. FL: UB Trier 55/BB/c12363.

[45] Hans Schupp. Ein (üb?)zeugendes problem. In C. Selter and G. Walther, editors, Mathematikdidaktik als design science: Festschrift für Erich Christian Wittmann, pages 188–195. Ernst Klett Grundschulverlag, 1999. ISBN 3-12-200060-1. Ordnung von b in (Z)/(mb − 1)(Z).

[46] H. Sewerin. Problemlösen in Klausuren: Ohne Strategie – aber wie? Der Mathematikunterricht, 25(1):69–79, 1979. HB: Z5577.

[47] Edward Allen Silver, editor. Teaching and Learning Mathematical Problem Solving: Multiple Research Perspectives (San Diego 1983). Erlbaum, 1985. ISBN 0-89859-681-5, 0-89859-759-5. HBZ 290, 361, 464, 465, 466, 467, 468. Enthält [7].

[48] Alexander Soifer. Mathematics as Problem Solving. MAA, 1987. ISBN 0-940263-00-9. The book is a practical guide to problem solving in secondary geometry, algebra, number theory, and combinatorics. Approximately 200 problems selected for their beauty, challenge, and elegance are presented here.

[49] M. Stein and B. Burchartz. Problemlöseprozesse bei Aufgaben mit nicht erreichbarer Zielsetzung – Methodische Probleme der qualitativen und quantitativen Auswertung umfangreichen Datenmaterials über Problemlöseprozesse. mathematica didactica, 21(1):21–44, 1998.

[50] James W. Stigler, Clea Fernandez, and Makoto Yoshida. Traditions of school mathematics in Japanese and American elementary classrooms. In L. P. Steffe et al., editor, Theories of Mathematical Learning, pages 149–175. Erlbaum, 1996. FL: UB Duisburg.

[51] James Tanton. Solve This. Math Activities for Students & Clubs. Classroom Resource Materials. MAA, 2001. ISBN 0-88385-717-0. Suitable for use as a liberal arts mathematics course in college or secondary school. This book irresistibly tempts the reader to embark on a journey of investigation and discovery.

[52] Terence Tao. Solving Mathematical Problems. A Personal Perspective. Oxford Univ. Press, 2006. ISBN 0-19-920560-4 Paperback, 0-19-920561-2 Hardback. Leads the reader through the tactics of solving mathematical problems. This text is ideal for students and for general readers of 14 years and above in pure mathematics.

[53] Günter Törner. Problemlösen in der Sekundarstufe I - Bericht über eine Tesreihe. Beiträge zum Mathematikunterricht, 1989:370–373, 1989. HB: Bb1256-1989.

[54] Heinz Wambach, editor. Die Förderung mathematischer Begabungen in der Sekundarstufe I (Tagungsbericht mit Praxisbeispielen für Fördermöglichkeiten), Bildung und Begabung e. V. Verlag Karl Heinrich Bock, Bad Honnef, 1988. ISBN 3-87066-153-4. S. 54: Fragenkatalog zur Systematisierung des Probierverhaltens (aus SEWERIN, Mathematische Schülerwettbewerbe); S. 55: Ratschläge zur Heuristik (aus: POLYA, Schule des Denkens).

[55] H. Winter. Entdeckendes Lernen im Mathematikunterricht: Einblicke in die Ideengeschichte und ihre Bedeutung für die Pädagogik. Didaktik der Mathematik. Vieweg, 1989. HB: Kb5960. ISBN 3-528-08978-4. S. 178-179: Polyas Heuristik des Problemlösens: (1) Verstehen der Aufgabe, (2)Ausdenken eines Plans, (3) Ausführen des Plans, (4) Rückschau; jeweils mit Detailfragen.

[56] Melanie Matchett Wood. Art of Problem Solving: a new resource for outstanding mathematics students. Focus (MAA), 27(3):10, 2007. Vgl. [?].

[57] Paul Zeitz. The Art and Craft of Problem Solving. Wiley, 1999. ISBN 0-471-13571-2. HBZ. Hinweis in Review auf S. 329 in CMJ 37:4 (2006).